Question 5. (15 pts) Find the maximum and minimum of f(t,y) = ry? on the circle x2 + y2 = 1
Answer #1

0 Griven Hay) = x y² on the circle x²ty²1 Then, y²1-22 Ha, y) = x (1-92) = X-2² zg(a) (say). Then, g(o)> 1-3x? and g(x)= -6At (73,6), 9(a)=-6.4+) the so So, fra, y) has minimum at (-2,1%) At (-43, -5%), g(x)= -6.6-7- en so so, fa,y) has minimum a